11-21-2018 06:56 AM
The 20G link for HA2 between the two PA-5220 firewalls (Active-Passive cluster) does not work correctly. It is a direct link using single-mode fiber and 10G-LR optics with a length of approximately 550 meters.
After restarting any of the two firewalls, the HA2 link does not raise (in the Down state). We only managed to lift it after disconnecting and reconnecting the fiber patch on either of the two firewalls.
On other occasions, the HSCI port enters a flaping loop (UP / Down) continuously. It is solved by disconnecting the fiber.
We have certified the single mode link with 20G-LR working fine(included patch cords).
The firmware version of the fws is 8.0.12.
Its seems like a bug but i can not find any bug related to this in 8.0.12.